### v1

@@ Abstract
Artifacts are a common problem in physiological time-series data collected from intensive care units (ICU) and other settings. They affect the quality and reliability of clinical research and patient care. Manual annotation of artifacts is costly and time-consuming, rendering it impractical. Automated methods are desired. Here, we propose a novel unsupervised approach to detect artifacts in clinical-standard minute-by-minute resolution ICU data without any prior labeling or signal-specific knowledge. Our approach combines a variational autoencoder (VAE) and an isolation forest (iForest) model to learn features and identify anomalies in different types of vital signs, such as blood pressure, heart rate, and intracranial pressure. We evaluate our approach on a real-world ICU dataset and compare it with supervised models based on long short-term memory (LSTM) and XGBoost. We show that our approach achieves comparable sensitivity and generalizes well to an external dataset. We also visualize the latent space learned by the VAE and demonstrate its ability to disentangle clean and noisy samples. Our approach offers a promising solution for cleaning ICU data in clinical research and practice without the need for any labels whatsoever.
Index Terms:
Unsupervised learning, Autoencoders, Artifact detection, Intensive care unit data, Isolation forests.
I Introduction
@@
We aim to develop unsupervised ML methods for automated artifact detection in minute-by-minute ICU time series. Building upon a labeled real-world dataset, we initially explore supervised ML techniques known for their reliable performance with labeled data. Then, we introduce a novel completely unsupervised ML approach that exhibits comparable performance to the fully supervised one. We validate the generalizability of this approach on recordings from an external dataset. Our methods leverage concepts from LSTMs, variational autoencoders (VAE), and decision trees. In particular, our contributions are:
@@
    A completely unsupervised approach to detect artifacts in minute-by-minute ICU data samples based on VAE and isolation Forest model (VAE-IF) that does not require previous labelling of artifacts and exploits averaged latent sample representations for enhanced accuracy.
@@
    A supervised model based on LSTM with self-attention (LSTM-ATTn) as an alternative for artifact detection when labels are available in a training set.
@@
    Experiments with a real-world ICU dataset to compare the performance of the alternative models for artifact detection and visualise the latent representations by the unsupervised approach.
The manuscript is organised as follows. Section II covers literature review. Section III introduces our methodology and benchmark techniques. Section IV then presents the experiments and results. Finally, Sections V to VII discuss and conclude our study.
II Literature review
@@
It is worth considering techniques developed for detecting anomalies in other kinds of time series as well since this problem closely resembles artifact detection in physiological time series. Munir et al. [20] developed a CNN-based model with a predictor and anomaly detector module. The predictor module predicts the next timestamp sample, which is then fed into the anomaly detector module. The anomaly detector module computed the Euclidian distance between actual sample and predicted sample, which determined if a sample is an outlier based on a set threshold. The method could detect both point and contextual anomalies in periodic time-series.
@@
In this work, we develop a completely unsupervised approach utilizing $\beta$-VAE [21] and Isolation forest [22] to detect artifacts at sample level in minute-resolution vital signs data, including mean blood pressure (BPm), heart rate (HRT) and mean intracranial pressure (ICPm), aiming to completely eliminate the need for labeled data when building models for artifact detection. Unlike previous studies utilizing VAEs to detect artifacts/anomalies by first training on clean data, our method trains on raw data containing both clean samples and artifacts. We also develop a fully supervised classifier based on LSTM with attention as a proposed benchmark [23, 24]. We also implement isolation forest on its own and XGBoost as benchmarks. We perform an ablation study to demonstrate the significance of feature extraction with a VAE and the importance of the self-attention component in the LSTM classifier. We also train an LSTM classier without the self-attention component to demonstrate its significance. Additionally, we explore the latent space of real-world data containing artifacts. In contrast with previous techniques, we train our model on a dataset of all the mentioned vital signs mixed. We show that the model learns and generalizes the dynamics of artifacts across the different vital signs. We validate our method using gold-standard annotations by an experienced expert and demonstrate the generalizability of the method on the MIMIC-IV [25] dataset without retraining.
III Methodology
In this Section, we present a novel unsupervised approach for artifact detection in minute-by-minute physiological ICU recordings, addressing the challenge of limited labeled data. As highlighted in Section I, the scarcity of labeled data makes our approach particularly valuable. While we recognize the scarcity of labeled data, we first establish a performance benchmark using supervised models. Then, we present and evaluate our novel unsupervised approach, showcasing its generalizability to an external dataset. However, before delving into the details of the models, we describe the dataset, preprocessing, and problem formulation.
III-A KidsBrainIT dataset
@@
We have access to a research dataset from the KidsBrainIT project [26] that was collected prospectively and extracted for research purposes from Paeditric Intensive Care Units (PICU). We use the KidsBrainIT pilot dataset, which contains about 98 patients from 2 PICU in the UK and about 27 patients that were added later on, making up a total of 125. Each patient in the dataset has around 14 physiological signals that were collected at minute-by-minute intervals, a routine clinical setup in ICU. It also includes annotations on artifacts by an expert researcher. For our experiments, we used the BPm, ICPm, and HRT as these signals have the least amount of missing values across all patients recordings. Fig. 1 shows the distribution of artifacts and good signal for the selected vital signs.
@@ III-B Data preprocessing
III-C Problem formulation
We consider a preprocessed dataset of size $N$, $D:=\left\{\left(\textbf{x}_{i},\textbf{y}_{i}\right)\right\}$ such that $\textbf{x}\in\mathbb{R}^{W},\textbf{y}\in\left\{0,1\right\}^{W},\mathit{i}\in\left\{1\ldots N\right\}$, where $\textbf{x}_{i}$ is a segment of length $W$ of the time-series and $\textbf{y}_{i}$ is the corresponding label for each sample in the segment as described in Section III-B. Additionally, we define $\textbf{z}\in\mathbb{R}^{H}$, a vector of size $H$, as the latent representation of a sample $x$ in the segment x. The latent representation for a signal of length $L$ is represented by a matrix $\textbf{Z}\in\mathbb{R}^{L\times H}$.
III-D Proposed benchmarks
In this section, we describe the supervised models to be used as a performance benchmarks for the unsupervised approach. For this task, we build two models: a neural-network model based on LSTM with self-attention, and XGBoost.
III-D1 LSTM with self-attention (LSTM-ATTn)
$\mathbf{h}_{t}=LSTM(x_{t},\mathbf{h}_{t-1}),$ (1)
where $\mathbf{h}_{t}\in\mathbb{R}^{H}$ is the hidden state vector and $x_{i}\in\mathbb{R}$ is a sample in segment $\mathbf{x}$ as defined in III-C. A Bidirectional LSTM [27] has access to past and future information. Its hidden state has an effective size of $\mathbf{h}_{t}\in\mathbb{R}^{2H}$.
where ${\mathbf{h}}^{\prime}_{i}$ is the final hidden state and $W$ is the window size as defined in Section III-B. We also explore the effect of using a bidirectional LSTM during hyper-parameter tuning. Next, we add a dense layer with a sigmoid activation function to give a probability estimate ${y}^{\prime}_{i}$ of sample $x_{i}$ being an artifact. We train the model using Adam optimizer [28]. Fig 3 shows the architecture diagram of the LSTM-Attn classifier.
III-D2 XGBoost
XGBoost [29], or Extreme Gradient Boosting, is a decision tree ensemble based on CART (Classification and Regression Trees) models trained using the gradient boosting algorithm [30]. The algorithm trains additively by fitting new trees to the residuals. XGBoost is known for its excellent performance [16], and we use it as a second benchmark.
III-E Fully unsupervised approach: VAE-IF
In this section, we present our proposed unsupervised approach for artifact detection in minute-by-minute physiological recordings. Notably, this approach is particularly valuable as it does not rely on labeled data. At a high level, our approach involves pre-training a VAE [31] model on the preprocessed training dataset, which includes both noisy and clean segments, $\mathbf{x}_{i}$, as described in Sections III-B and III-C. Subsequently, we utilize the encoder of the trained VAE to extract features from a test signal. Once the features are extracted, we directly employ the Isolation Forest [22] anomaly detection algorithm to classify samples in the signal as normal or artifact (anomaly). In the following subsections, we give a detailed description of the main components of this approach.
III-E1 Training the VAE
@@
where $\theta,\phi$ are the parameters of the decoder $p(\mathbf{x}|\mathbf{z})$ and encoder $q(\mathbf{z}|\mathbf{x})$ and $\beta$ is a parameter that can be tuned to encourage the model to learn disentangled representations in the latent space. The encoder consists of two stacked LSTM layers, followed by attention and a dense layer. Similarly, the decoder has a comparable architecture, except for the absence of the attention block. The first three layers of the encoder are similar to that of the LSTM-Attn model in Fig 3. The full architecture for the VAE can be seen in Fig. 4.
We train the VAE with a mix of clean and artifact-containing data segments x. We use Bayesian optimization (BO) [32] to tune model hyper-parameters listed in Table I.
Figure 3: Architectual diagram of the supervised LSTM-Attn benchmark
Figure 4: Architectural diagram of the VAE
III-E2 Feature extraction and averaging
@@
III-E3 Artifact detection with iForest
Isolation Forest (iForest) [22] is an algorithm for anomaly detection which uses an ensemble of isolation trees to partition data based on the assumption that anomalies exhibit distinct characteristics that set them apart from normal data points. Consequently, anomalies are expected to be isolated closer to the root node of the trees compared to normal data points. Compared to other algorithms like OCSVM, iForest is computationally faster and scales well to high-dimensional datasets [22].
We use Sklearn’s [33] implementation of iForest with its default hyperparameters for artifact detection. iForest produce a sequence with output values corresponding to clean and artifactual samples, represented by the vector sequence $\bar{Z}$. We first train the iForest model on the extracted features of all patients and the corresponding target signals in training set before using it on the testing set.
TABLE I: Hyperparameters and ranges included in the BO
@@ III-E3 Artifact detection with iForest
(b) LSTM-ATTn
(c) XGBoost
III-F Ablation study
@@
III-F1 Feature extraction ablation
@@
III-F2 Self-Attention Ablation
@@
IV Experiments and Results
We used BO to tune model hyperparameters for the VAE part of the VAE-IF, LSTM-ATTn, and XGBoost, using Geometric Mean as the objective function. We also explored training hyperparameters such as window size ($W$), learning rate, and number of epochs. The range of hyperparameters explored for the three models is summarised in Table I.
Using the test set, we evaluate the performance of VAE-IF and compare it to the supervised models; LSTM-ATTn and XGBoost. To measure performance, we compute sensitivity and specificity on a sample-by-sample basis. We compute the performance metrics from each recording in the test set separately and report the mean and standard deviation scored by each model.
@@ IV-A Unsupervised approach
IV-B Benchmarks
Similarly, we tune the hyperparameters for the LSTM-ATTn and XGBoost using BO. The best hyperparameters for the benchmark models also included in the supplementary material. Table II and II show the best performance on test set for LSTM-ATTn and XGBoost, respectively. Results show that LSTM-ATTn achieve the highest sensitivity for HRT signal. The benchmark models achieved higher specificity compared to VAE-IF across all three signals.
IV-C External validation
We tested the generalizability of VAE-IF model on sample patients from the MIMIC-IV [25] waveform dataset. Here, we present results for the ICP signal of patient p15649186. Given the difference in sampling rates between the datasets, we downsampled the ICP signal from 250Hz to 0.0167Hz (one sample per minute). We then applied the VAE-IF without any retraining. Fig. 6 shows the model’s artifact prediction and the patient recording. Additional examples are included in the supplementary material.

### v2

@@ Abstract
Artifacts are a common problem in physiological time series collected from intensive care units (ICU) and other settings. They affect the quality and reliability of clinical research and patient care. Manual annotation of artifacts is costly and time-consuming, rendering it impractical. Automated methods are desired. Here, we propose a novel fully unsupervised approach to detect artifacts in clinical-standard, minute-by-minute resolution ICU data without any prior labeling or signal-specific knowledge. Our approach combines a variational autoencoder (VAE) and an isolation forest (IF) into a hybrid model to learn features and identify anomalies in different types of vital signs, such as blood pressure, heart rate, and intracranial pressure. We evaluate our approach on a real-world ICU dataset and compare it with supervised benchmark models based on long short-term memory (LSTM) and XGBoost and statistical methods such as ARIMA. We show that our unsupervised approach achieves comparable sensitivity to fully supervised methods and generalizes well to an external dataset. We also visualize the latent space learned by the VAE and demonstrate its ability to disentangle clean and noisy samples. Our approach offers a promising solution for cleaning ICU data in clinical research and practice without the need for any labels whatsoever.
keywords:
Unsupervised learning , Autoencoders , Artifact detection , Intensive Care Unit data , Isolation Forest , time series
[…]
[label1]organization=Institute for Imaging, Data and Communications (IDCOM), School of Engineering, University of Edinburgh,city=Edinburgh, postcode=EH9 3FB, country=U.K.
[…]
[label5]organization=Corresponding author,addressline=hhaule@ed.ac.uk
1 Introduction
@@
We aim to develop unsupervised ML methods for automated artifact detection in minute-by-minute ICU time series. Building upon a labeled real-world dataset, we are able to implement supervised ML techniques known for their reliable performance with labeled data as benchmarks. Importantly, we introduce a novel completely unsupervised ML approach that exhibits comparable performance to the fully supervised benchmarks. We validate the generalizability of this approach on recordings from an external dataset. Our methods leverage concepts from LSTMs, variational autoencoders (VAE), and decision trees. In particular, our contributions are:
@@
    A completely unsupervised hybrid model to detect artifacts in routinely collected ICU data samples based on VAE and isolation Forest model (VAE-IF) that does not require any previous labeling of artifacts and exploits averaged latent sample representations for enhanced accuracy.
@@
    Implementations of supervised benchmarking models (including deep learning and traditional statistical methods) and comparisons of the performance of our VAE-IF model to that of the fully supervised models for artifact detection in ICU data.
@@
    Experiments with real world data to explore the learned latent space of VAE-IF, demonstrating its ability to disentangle artifacts from clean data despite being unsupervised.
The manuscript is organised as follows. Section 2 covers literature review. Section 3 introduces our methodology and benchmark techniques. Section 4 then presents the experiments and results. Finally, Sections 5 to 6 discuss and conclude our study.
2 Literature review
@@
It is worth considering techniques developed for detecting anomalies in other kinds of time series as well since this problem closely resembles artifact detection in physiological time series. Munir et al. [20] developed a CNN-based model with a predictor and anomaly detector module. The predictor module predicts the next timestamp sample, which is then fed into the anomaly detector module. The anomaly detector module computed the Euclidian distance between actual sample and predicted sample, which determined if a sample is an outlier based on a set threshold. The method could detect both point and contextual anomalies in periodic time series.
Numerous studies have used VAEs for artifact detection [17] and anomaly detection [21, 22, 23], including a $\beta$-VAE [24]. However, all of these studies train their VAEs with data that do not contain anomalies, making all these approaches “weakly” unsupervised since labels are indeed required to filter the data prior to training, as remarked by Gangloff et al. [23]. In contrast, our proposed approach is “completely” unsupervised, since we train our VAE on unfiltered data. For training their models, Lin et al. and Niu et al. split a time series recording into two parts: the first part for training and the last for testing. This strategy requires a model to have seen part of the time series and is less practical for detecting artifacts/anomalies on retrospective data. Our approach splits whole recordings into training and testing sets, ensuring no information leakage from the training to the test set.
3 Methodology
@@
In this work, we develop a completely unsupervised approach utilizing $\beta$-VAE [25] and IF [26] to detect artifacts at the sample level in minute-resolution vital signs data, including mean blood pressure (BPm), heart rate (HRT), and mean intracranial pressure (ICPm). This approach aims to eliminate the need for labeled data when building models for artifact detection. Unlike previous studies utilizing VAEs to detect artifacts/anomalies by first training on clean data [24, 17, 21, 22, 23], our method trains on raw data containing both clean samples and artifacts. We show that the model learns and generalizes the dynamics of artifacts across different vital signs by training our model on a dataset of all the mentioned vital signs mixed. Additionally, we develop a fully supervised classifier based on LSTM with attention as a proposed benchmark [27, 28], alongside implementing Isolation Forest on its own, XGBoost and ARIMA as additional benchmarks. An ablation study demonstrates the significance of feature extraction with a VAE and the importance of the self-attention component in the LSTM classifier, including training an LSTM classifier without the self-attention component to highlight its significance. We also explore the latent space of real-world data containing artifacts. We validate our method using gold-standard annotations by an experienced expert and demonstrate its generalizability on the MIMIC-IV [29] dataset without retraining. As highlighted in Section 1, the scarcity of labeled data makes our approach particularly valuable. While recognizing this scarcity, we first establish a performance benchmark using supervised models before presenting and evaluating our novel unsupervised approach, showcasing its generalizability to an external dataset. Before delving into the details of the models, we describe the dataset, preprocessing, and problem formulation.
3.1 KidsBrainIT dataset
@@
We have access to a research dataset from the KidsBrainIT project [30] that was collected prospectively and extracted for research purposes from Paeditric Intensive Care Units (PICU) with approval reference number 17/WS/0086 by the Research Ethics Committee in the United Kingdom. We use the KidsBrainIT pilot dataset, which contains about 98 patients from 2 PICU in the UK and about 27 patients that were added later on, making up a total of 125. Each patient in the dataset has around 14 physiological signals that were collected at minute-by-minute intervals, a routine clinical setup in ICU. It also includes annotations on artifacts by an expert researcher. For our experiments, we used the BPm, ICPm, and HRT as these signals have the least amount of missing values across all patients recordings. Fig. 1 shows the distribution of artifacts and good signal for the selected vital signs.
@@ 3.2 Data preprocessing
Figure 3: Figure showing concatenated signals across all patients showing that the three signals are non-periodic
Mean Std Dev Variance Min Max
BPm 85.53 12.46 155.28 23.0 215.0
ICPm 20.30 13.77 189.65 -18.0 158.0
HRT 98.35 20.86 435.07 0.0 338.0
BPm ICPm HRT
BPm 1.00 0.21 -0.04
ICPm 0.21 1.00 0.06
HRT -0.04 0.06 1.00
Table 1: Statistical properties and correlation matrix showing weak or no dependency between BPm, ICPm and HRT
3.3 Problem formulation
We consider a preprocessed dataset of size $N$, $D:=\left\{\left(\textbf{x}_{i},\textbf{y}_{i}\right)\right\}$ such that $\textbf{x}\in\mathbb{R}^{W},\textbf{y}\in\left\{0,1\right\}^{W},\mathit{i}\in\left\{1\ldots N\right\}$, where $\textbf{x}_{i}$ is a segment of length $W$ of the time series and $\textbf{y}_{i}$ is the corresponding label for each sample in the segment as described in Section 3.2. Additionally, we define $\textbf{z}\in\mathbb{R}^{H}$, a vector of size $H$, as the latent representation of a sample $x$ in the segment x. The latent representation for a signal of length $L$ is represented by a matrix $\textbf{Z}\in\mathbb{R}^{L\times H}$.
3.4 Fully unsupervised approach: VAE-IF
In this section, we present our proposed unsupervised approach for artifact detection in minute-by-minute physiological recordings. Notably, this approach is particularly valuable as it does not rely on labeled data. At a high level, our approach involves pre-training a VAE [31] model on the preprocessed training dataset, which includes both noisy and clean segments, $\mathbf{x}_{i}$, as described in Sections 3.2 and 3.3. Subsequently, we utilize the encoder of the trained VAE to extract features from a test signal. Once the features are extracted, we directly employ the IF [26] anomaly detection algorithm to classify samples in the signal as normal or artifact (anomaly). In the following subsections, we give a detailed description of the main components of this approach.
3.4.1 Training the VAE
@@
where $\theta,\phi$ are the parameters of the decoder $p(\mathbf{x}|\mathbf{z})$ and encoder $q(\mathbf{z}|\mathbf{x})$ and $\beta$ is a regularization coefficient. Tuning $\beta$ has a smoothing effect on the high-frequency components of the signals, enabling the model to learn stable representations, which is our main motivation for using this variant of VAEs. The encoder consists of two stacked LSTM layers, followed by attention and a dense layer. Similarly, the decoder has a comparable architecture, except for the absence of the attention block. The first three layers of the encoder are similar to that of the LSTM-Attn benchmark model in Fig 4. The operation of these LSTM blocks is described in Section 3.5.1. The full architecture for the VAE can be seen in Fig. 5.
We train the VAE with a mix of clean and artifact-containing data segments x. We use Bayesian optimization (BO) [32] to tune model hyper-parameters listed in Table 6(a).
Figure 4: Architectual diagram of the supervised LSTM-Attn benchmark.
Figure 5: Architectural diagram of the VAE.
3.4.2 Feature extraction and averaging
@@
3.4.3 Artifact detection with IF
IF [26] is an algorithm for anomaly detection which uses an ensemble of isolation trees to partition data based on the assumption that anomalies exhibit distinct characteristics that set them apart from normal data points. Consequently, anomalies are expected to be isolated closer to the root node of the trees compared to normal data points. Compared to other algorithms like OCSVM, IF is computationally faster and scales well to high-dimensional datasets [26].
We use Sklearn’s [33] implementation of IF with its default hyperparameters for artifact detection. IF produce a sequence with output values corresponding to clean and artifactual samples, represented by the vector sequence $\bar{Z}$. We first train the IF model on the extracted features of all patients and the corresponding target signals in training set before using it on the testing set.
Table 2: Hyperparameters and ranges included in the BO.
@@ 3.4.3 Artifact detection with IF
3.5 Proposed benchmarks
In this section, we describe the supervised models to be used as a performance benchmarks for the unsupervised approach. For this task, we: build a neural-network model based on LSTM with self-attention, and also use XGBoost and ARIMA, a standard statistical model.
3.5.1 LSTM with self-attention (LSTM-ATTn)
$\mathbf{h}_{t}=LSTM(x_{t},\mathbf{h}_{t-1}),$ (2)
where $\mathbf{h}_{t}\in\mathbb{R}^{H}$ is the hidden state vector and $x_{i}\in\mathbb{R}$ is a sample in segment $\mathbf{x}$ as defined in 3.3. A Bidirectional LSTM [34] has access to past and future information. Its hidden state has an effective size of $\mathbf{h}_{t}\in\mathbb{R}^{2H}$.
where ${\mathbf{h}}^{\prime}_{i}$ is the final hidden state and $W$ is the window size as defined in Section 3.2. We also explore the effect of using a bidirectional LSTM during hyper-parameter tuning. Next, we add a dense layer with a sigmoid activation function to give a probability estimate ${y}^{\prime}_{i}$ of sample $x_{i}$ being an artifact. We train the model using Adam optimizer [35]. Fig 4 shows the architecture diagram of the LSTM-Attn classifier.
3.5.2 XGBoost
XGBoost [36], or Extreme Gradient Boosting, is a decision tree ensemble based on CART (Classification and Regression Trees) models trained using the gradient boosting algorithm [37]. The algorithm trains additively by fitting new trees to the residuals. XGBoost is known for its excellent performance [16], and we use it as a second benchmark. We use XGBoost to classify each part of a signal as either an artifact or not. The input is the original signal, and the output is a label for each sample in that signal.
3.5.3 ARIMA
In addition to ML-based models, we benchmark using an ARIMA [38] model, which is a popular statistical model for time series analysis. Autoregressive Integrated Moving Average (ARIMA) is a statistical model for representing a stochastic non-stationary signal, assuming the signal has stationary and non-stationary components. The non-stationary component is represented by stochastic autoregressive (AR) process of order $P$ (i.e. the number of steps in the past required to predict the current value). The stationary component is represented by a moving average (MA) process of order $Q$ (i.e. the number of neighboring steps to average). To make the signal stationary, differencing operations of order $D$ are applied to the signal. We fit a signal with ARIMA model and treat residuals more than three standard deviations from the mean as artifacts.
3.6 Ablation study
@@
3.6.1 Feature extraction ablation
@@
3.6.2 Self-Attention Ablation
@@
4 Experiments and Results
We used BO to tune model hyperparameters for the VAE part of the VAE-IF, LSTM-ATTn, and XGBoost, using Geometric Mean as the objective function. We also explored training hyperparameters such as window size ($W$), learning rate, and number of epochs. We use the function “auto_arima” from the package pmdarima[39] to find the best ARIMA model (i.e., $P$, $Q$, $D$) for each of the three signals in the training set. The range of hyperparameters explored for the other models is summarised in Table 2.
Using the test set, we evaluate the performance of VAE-IF and compare it to ARIMA and the supervised machine models: LSTM-ATTn and XGBoost. To measure performance, we compute sensitivity and specificity on a sample-by-sample basis. We compute the performance metrics from each recording in the test set separately and report the mean and standard deviation scored by each model.
@@ 4.1 Unsupervised approach
4.2 Benchmarks
Similarly, we tune the hyperparameters for the LSTM-ATTn and XGBoost using BO. The best hyperparameters for the benchmark models also included in the supplementary material. Table 3 show the best performance on test set for LSTM-ATTn, XGBoost, and ARIMA. Results show that LSTM-ATTn achieve the highest sensitivity for HRT signal. The benchmark models achieved higher specificity compared to VAE-IF across all three signals.
4.3 External validation
We tested the generalizability of VAE-IF model on sample patients from the MIMIC-IV [29] waveform dataset. Here, we present results for the ICP signal of patient p15649186. Given the difference in sampling rates between the datasets, we downsampled the ICP signal from 250Hz to 0.0167Hz (one sample per minute). We then applied the VAE-IF without any retraining. Fig. 7 shows the model’s artifact prediction and the patient recording. Additional examples from more than 60 other subjects are included in the supplementary material.